# Examining the Prevalence and Dynamics of AI-Generated Media in Art Subreddits

### v1

@@ Abstract
Broadly accessible generative AI models like Dall-E have made it possible for anyone to create compelling visual art. In online communities, the introduction of AI-generated content (AIGC) may impact community dynamics by shifting the kinds of content being posted or the responses to content suspected of being generated by AI. We take steps towards examining the potential impact of AIGC on art-related communities on Reddit. We distinguish between communities that disallow AI content and those without a direct policy. We look at image-based posts made to these communities that are transparently created by AI, or comments in these communities that suspect authors of using generative AI. We find that AI posts (and accusations) have played a very small part in these communities through the end of 2023, accounting for fewer than 0.2% of the image-based posts. Even as the absolute number of author-labelled AI posts dwindles over time, accusations of AI use remain more persistent. We show that AI content is more readily used by newcomers and may help increase participation if it aligns with community rules. However, the tone of comments suspecting AI use by others have become more negative over time, especially in communities that do not have explicit rules about AI. Overall, the results show the changing norms and interactions around AIGC in online communities designated for creativity.
Introduction
@@
Online communities are grappling with how to respond to the availability of AI-generated content (AIGC) online. In this paper, we seek to characterize how online visual art communities may diverge based on whether they expressly forbid AI or choose to remain neutral towards the use of AI. On the one hand, communities that forbid the use of AI often cite that content quality may be lower, creative perspectives will be homogenized, and community values could be compromised (Lloyd, Reagle, and Naaman 2023; Frosio 2023; Radivojevic et al. 2024). Conversely, communities that choose not to forbid AI may be calling upon arguments that AI can be used in ways that help democratize certain hobbies, spark creativity, and may increase participation (Gero, Long, and Chilton 2023; Lloyd, Reagle, and Naaman 2023; Panchanadikar et al. 2024). Online communities are wrestling with these choices and strategies. Characterizing the dynamics and behaviors of community members, especially in communities that differ in their rules toward AI, can inform future community actions. We explore how this tension in explicitly restricting AI is playing out within art subreddit communities, which are highly creative, have specific community values, and where members may have broad access to AI for image generation.
@@
More broadly, we also provide an initial understanding of behaviors around AIGC that have developed in art-based online communities. AIGC has been described as a “paradigm shift in content creation and knowledge representation” (Wang et al. 2023). Across different fields, emergent work has attempted to quantify the prevalence and responses to AIGC online, from impact to academic peer review and synthetic news articles, to online crowdwork  (Liang et al. 2024; Latona et al. 2024; Veselovsky, Ribeiro, and West 2023; Christoforou, Demartini, and Otterbacher 2024; Hanley and Durumeric 2024). However, there is a need to understand how AIGC may impact online social ecosystems, since these are often based on community values and authenticity in ways that may be called into question by the widespread adoption of AIGC (Lloyd, Reagle, and Naaman 2023). In this study, we tackle this question as it related to online art communities.
@@
To this end, we collect a list of 57 art-related subreddit communities, which we label as AI-neutral or AI-disallowed based on their stance towards AI posts. From these subreddits, we retrieve and classify 28,405 image-based posts where the poster has either admitted to generating the image with AI (Transparent AI post), or a commenter has accused the image of being AI-generated (Suspected AI post). We then analyze the number of AI-generated posts made to these communities over time, the trajectories of newcomers to these communities who post AI-generated images, and the speed and impact of AI-suspecting comments on posts.
From our analysis of this data, we make the following contributions:
  • Overall, we find that the prevalence of posts made to these art-based communities where the authors transparently self-report AI-generated use, or where commenters raise suspicion about the use of AI is surprisingly low, affecting less than 0.2% of image-based posts in our dataset.
@@
  • We find that the production of AI-based image posts started increasing in early 2022, peaked in August 2022 and has decreased since, while comments suspecting the use of AI started rising only months later, and may have not declined as much. These trends were similar between communities that disallow AI and those who do not explicitly do so.
  • We identify that AI posts are more likely to be used by newcomers to a community, and that newcomers who initially share AI-generated images often continue to share AI-generated images, indicating that their interest in these communities has been enabled by the availability of these technologies.
  • We find that image-based posts that are suspected in using AI have been receiving exceedingly negative comments over the period of the data collection.
@@
Overall, our findings set the stage for understanding the prevalence, impact, and reception of AIGC in online visual art communities.
Background
@@
Online communities have been a popular topic of study for internet researchers. One relevant theory to online communities is the concept of communities of practice (CoP) (Wenger 1999, 1996; Lave 1991; McDonald and Cater-Steel 2016; Wenger, McDermott, and Snyder 2002), where groups of people who share a common interest come together to explore these ideas, share, and improve their practice. CoP theory emphasizes that members engage in collective learning through shared practices, discussions, and activities, as is the case in Reddit art-based communities. Broadly speaking, online communities often wrestle with tradeoffs between different priorities, such as growth and attracting newcomers, sustainability, content quality, and fostering a sense of virtual belonging (Kairam, Mercado, and Sumner 2022; Lin et al. 2017; Hwang and Foote 2021; Kraut et al. 2012). Recent work on Reddit mapped subreddits into five archetype, which relate to different ways they produce a sense of community (Prinster et al. 2024) In this work, we are interested in art communities, which are probably most aligned with the Content Generation archetype from that work, and aim to understand how such a community of practice might be impacted by AIGC.
@@
Indeed, the recent emergence of AIGC may present a boon for online communities, but could also put community values under threat. The Diffusion of Innovations theory, formulated by Rogers (Rogers, Singhal, and Quinlan 2014; Rogers and Adhikarya 1979), focuses on the process by which individuals adopt new innovations and ideas, suggesting that different communities may also adapt innovations in diverse ways.  Weld, Zhang, and Althoff (2024) identify that Reddit communities put emphasis on diverse values, and often prioritizing one comes at the cost of another – for example, the quality of content can be in direct conflict with the need to grow communities and be self-sustaining. Prior work has identified that the choice for whether to allow or disallow AI can force subreddit moderators to contend with which of these values to codify as rules (Lloyd, Reagle, and Naaman 2023). These decisions can have consequences beyond individual subreddits; recent work has demonstrated that simple changes like exposing moderation decisions significantly impacts the behaviors of users on the Reddit platform (Jhaver, Rathi, and Saha 2024). As communities content with the decision of how to moderate AI-generated content, this study provides initial evidence for how these decision may shape participation in online communities.
AI-Generated Images
@@
Text-to-image models, such as DALL-E [^1] https://openai.com/index/dall-e-3/, Stable Diffusion [^2] https://stability.ai/, and Midjourney [^3] https://www.midjourney.com, have exhibited exceptional performance in generating both abstract and photorealistic images. Prior to the advent of these models, Generative Adversarial Networks (GAN)-based models were employed to generate images (Brock, Donahue, and Simonyan 2018; Goodfellow et al. 2014; Zhang et al. 2021). With such popular image-generation tools available since 2022, the concept of AI-generated content have been increasingly impacting online communities.
@@
Since these tools are so advanced, one key challenge is the automatic detection or manual identification of AI-generated images. Efforts to determine whether an image was generated by generative models reveal the challenges humans face in accurately identifying AI-generated images (Lu et al. 2024; Nightingale and Farid 2022), with a misclassification rate of 38.7% (Lu et al. 2024). Furthermore, even the top-performing AI-generated image detection model achieves a misclassification rate of 13% (Lu et al. 2024).
@@
A few papers have specifically examined the detection of AI-generated art (Ha et al. 2024; Wang, Huang, and Hong 2023).  Ha et al. (2024) showed that an approach combining human experts and automatic detectors was most effective, while non-artist crowdworkers generally were unable to discern between human-created art and AI-generated images.
@@
AI-generated Art and Online Communities
@@
AI-generated art is not just as an attempt for fabrication or low-effort creation. Several papers make the case for AI-generated art creations as authentic and creative (Mazzone and Elgammal 2019; Garcia 2024).  Garcia (2024) specifically investigates AI-generated art produced by advanced technologies like DALL-E [^4] https://openai.com/index/dall-e-3/. The paper provides a comprehensive analysis of the challenges associated with generative AI in art, including issues of authenticity, intellectual property, ethical concerns, the influence on traditional art practices, and market oversaturation.
@@
The discussions of challenges and opportunities in AI art extend to the context of online art communities. Communities of practice may have reasons to adopt or reject AI-generated art. For example, AI art may be seen as not contributing to improving their practice, or lacking creativity. On the other hand, AI art could be viewed as a new avenue for contribution, or a way to lower the barriers for newcomers to participate (Kraut et al. 2012).
@@
Some recent works directly explore the impact of AI-generated images on different artistic communities  (Wei and Tyson 2024; Guo et al. 2024; Zhou and Lee 2024). Wei and Tyson (2024) examined the spread of AI-generated images on Pixiv, a Japanese online community for artists, where artists are required to explicitly label their submissions as either human-generated or AI-generated. The paper reveals that the introduction of AI led to a 50% increase in new artworks, but no corresponding increase in the number of views or comments. Additionally, the introduction of AI prompted an increase in the number of new creators, but a 4.3% decrease in new creators of human-generated content. Another exploratory work (Guo et al. 2024) focused on the artist community DeviantArt which also explicitly allows AIGC. Despite DeviantArt’s requirement for users to indicate the use of AI techniques, the authors identified AI-generated artworks that did not explicitly disclose AI usage, exposing the complicated motivation of contributors. Zhou and Lee (2024) looked at data from an undisclosed online community focused on the creativity and productivity of artists who adopted generative AI, providing evidence for a broadening of participation and success in the community examined.
@@
The papers mentioned above examined dedicated art communities, with centralized policies and rules. In this work, we are focused on Reddit, and particularly on the decentralized, mostly-self governed (Fiesler et al. 2018) communities on this site, and especially those dedicated to art. Recent work has shown that subreddit moderators are already grappling with AIGC (Lloyd, Reagle, and Naaman 2023). Subreddit moderators are creating rules that regulate AI use in their communities, clarify norms, and enforce policies. By leveraging the organizational structure of Reddit’s subreddits, we are able to analyze data based on the distinct characteristics of different communities, including their own rules — namely, if the community explicitly disallows AI, or is neutral towards it in its rules. Our analysis also covers questions such as submissions volume and number of participants, in light of the different community types.
Dataset
@@
To investigate online discussions within art communities regarding AI-generated images, we undertook a careful process to select data from Reddit communities (subreddits) dedicated to art. Then we used an initial keyword-based filtering using a list of generative AI keywords, presented in this section, and LLM-based labeling, detailed in the next section.
@@
Since there is no known existing list of art-related subreddits from prior work, we used the Reddit API combined with several external sources to construct an initial comprehensive list. These sources include art-related subreddits retrieved using the PRAW API’s search function [^5] https://praw.readthedocs.io/en/latest/index.html, an article detailing Reddit’s largest communities [^6] https://www.reddit.com/best/communities/1/, and other unofficial articles describing the best art subreddits [^7] https://www.reddit.com/r/redditlists/comments/141nga/list˙of˙art˙subreddits/, https://csslab.cs.toronto.edu/gs/#interactive. This initial step yielded a list of 276 subreddits. We further narrowed this initial list of subreddits by manually examining the description of each subreddit to ascertain its relevance to art. When descriptions were insufficient for a decision, we examined a random sample of five posts from the subreddit. A subreddit was considered to be related to art if it featured visual art, regardless of whether the art posted in the subreddit is created by the authors of the posts, and could be focused on a specific genre of art, or be more general. Subreddits dedicated to memes, videos, or nudity were excluded. This elimination step resulted in 141 relevant subreddits. Finally, since we wanted to focus on AI use and response to AI in broader art communities, we excluded the “aiArt” subreddit, resulting in 140 relevant subreddits.
@@
After identifying the relevant subreddits, we proceeded to collect corresponding posts. We utilized publicly available data hosted on Academic Torrents, which contains files for the top 40,000 subreddits, spanning from the beginning through the end of 2023 (Watchful1 2024). This data is mostly building on the Pushshift API (Baumgartner et al. 2020). Each data record corresponds to the Pushshift structure [^8] https://github.com/pushshift/api, including fields such as ID, date, title, author, and attached media. Given that these torrents only cover the top 40,000 subreddits, we successfully obtained torrents for 57 subreddits from our list, resulting in a final list of 57 subreddits for our analysis (see Appendix A).
@@
With the post and comment data from the subreddit submissions, we set out to identify potential discussions about generative AI by filtering for submissions that include media items, and by using a list of AI-related keywords. First, we focused only on posts that included a media item, such as an image or video. This focus allows us to analyze discussions related to specific AI-generated art instances, rather than general discussions about art and generative AI. Second, we filtered these posts using a list of keywords likely to appear in discussions about AI-generated images. We began with an initial list of keywords and searched for them within our set of subreddits. This initial list included various keywords, starting from general phrases such as “generative art” and ending with specific text-to-image tools names. By examining the discussions in the retrieved posts, we iteratively refined, filtered, and expanded the list of keywords. Each of these keywords might be a word or a phrase, and mostly includes combinations with the phrases “AI” or “artificial intelligence”, or names of text-to-image tools for creating AI-generated images. This keywords selection process was done iteratively and purposefully to achieve a robust set, resulting in a final set of 68 keywords (see Appendix B). We then identified occurrences of these 68 keywords in the titles, bodies, and comments of the posts in our 57 subreddits. This search yielded 28,864 posts or comments containing at least one of the keywords.
@@
Our final dataset includes 28,405 text items related to AI-generated images, from 14,574 distinct posts. Of these, 6,393 (43.9%) posts feature the keywords in the title or body, while 8,181 (56.1%) posts have keywords only in the comments. The dataset includes several fields: ID, post title, post body, complete comments tree, author, timestamp, detected keywords, subreddit name, and more. In the next sections, we enhance the dataset with the classification of the texts based on their relatedness to claims about AI-generative images, as well as the categorization of the included subreddits. The dataset, with the classification labels and subreddit categories, will be made available upon publication.
Classification of the Dataset Using LLMs

### v2

@@ Abstract.
Broadly accessible generative AI models like Dall-E have made it possible for anyone to create compelling visual art. In online communities, the introduction of AI-generated content (AIGC) may impact social dynamics, for example causing changes in who is posting content, or shifting the norms or the discussions around the posted content if posts are suspected of being generated by AI. We take steps towards examining the potential impact of AIGC on art-related communities on Reddit. We distinguish between communities that disallow AI content and those without such a direct policy. We look at image-based posts in these communities where the author transparently shares that the image was created by AI, and at comments in these communities that suspect or accuse authors of using generative AI. We find that AI posts (and accusations) have played a surprisingly small part in these communities through the end of 2023, accounting for fewer than 0.5% of the image-based posts. However, even as the absolute number of author-labeled AI posts dwindles over time, accusations of AI use remain more persistent. We show that AI content is more readily used by newcomers and may help increase participation if it aligns with community rules. However, the tone of comments suspecting AI use by others has become more negative over time, especially in communities that do not have explicit rules about AI. Overall, the results show the changing norms and interactions around AIGC in online communities designated for creativity.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The ACM Conference on Computer-Supported Cooperative Work and Social Computing; ; isbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Online communities are grappling with how to respond to the availability of AI-generated content (AIGC) online. On the one hand, communities that forbid the use of AI often cite that content quality may be lower, creative perspectives will be homogenized, and community values could be compromised (Lloyd et al., 2025b; Frosio, 2023; Radivojevic et al., 2024; Lloyd et al., 2025a). Conversely, communities that choose not to forbid AI may be calling upon arguments that AI can be used in ways that help democratize certain hobbies, spark creativity, and may increase participation (Gero et al., 2023; Lloyd et al., 2025b; Panchanadikar et al., 2024). As online communities wrestle with these choices and strategies, characterizing the dynamics and behaviors of community members, especially in communities that differ in their rules toward AI, can inform future community actions. We explore how this tension in explicitly restricting AI plays out within art subreddit communities, which are highly creative, have specific community values, and where members may have broad access to AI for image generation.
@@
AIGC has been described as a “paradigm shift in content creation and knowledge representation” (Wang et al., 2023b), and already has significant impact on online information. Across different fields, emergent work has attempted to quantify the prevalence and responses to AIGC online, from impact to academic peer review and synthetic news articles, to online crowdwork  (Liang et al., 2024; Latona et al., 2024; Veselovsky et al., 2023; Christoforou et al., 2024; Hanley and Durumeric, 2024; Vykopal et al., 2023). However, there is a need to understand how AIGC may impact online social ecosystems, since these are often based on community values and authenticity in ways that may be called into question by the widespread adoption of AIGC (Lloyd et al., 2025b, a). In this study, we tackle this question as it relates to online art communities.
We are interested in how generative AI changes some core aspects of online community life. The first aspect is content production. Online communities lean on the active production of relevant content through posts and discussions to fulfill their purpose and maintain their perceived popularity (Gero et al., 2023; Panchanadikar et al., 2024; Atcheson et al., 2024). With generative AI, we may expect shifts in content production, both in terms of quantity and perceived quality, as AI tools change who can create and what they create (Wei and Tyson, 2024; Zhou and Lee, 2024). The second aspect is participation. Online communities rely on participation, and in this regard, practices, policies, and norms around participation barriers often dictate the dynamics and even success of communities (Atcheson et al., 2024; Frosio, 2023; Radivojevic et al., 2024; Lloyd et al., 2025a). AIGC may influence who participates, potentially lowering barriers for newcomers or conversely, deterring some contributors (Kraut et al., 2012; Wei and Tyson, 2024). The third aspect is the communication and trust between community members. In online communities, it is often critical that members trust each other and have a positive exchange (Mazzone and Elgammal, 2019). AIGC may alter how members engage with each other and the overall social dynamics in the community (Guo et al., 2024; Garcia, 2024). We therefore address the following research questions:
  • RQ1: How prevalent are AIGC posts within online art communities, and how does the volume of such posts evolve over time?
  • RQ2: Does AIGC facilitate the participation of new creators in these communities, and do they remain active contributors over time?
  • RQ3: How do community members react to and engage with posts containing media suspected of being AI-generated?
Moreover, we examine these research questions in the context of the communities’ approach towards generative AI. We examine whether these AI-related participation and community dynamics differ between art communities that forbid AIGC and those that remain neutral towards AI-generated content.
@@
To this end, we collect a list of 57 art-related subreddit communities, which we label as AI-neutral or AI-disallowed based on their stance towards AI posts. Having all posts in these subreddits, spanning from the beginning of each subreddit through the end of 2023 (Watchful1, 2024), we retrieve and classify 28,405 image-based posts where the poster has either admitted to generating the image with AI (which we call a Transparent AI post), or a commenter has accused the image of being AI-generated (Suspected AI post). We then analyze the number of AI-generated posts made to these communities over time, the trajectories of newcomers to these communities who post AI-generated images, and the speed and impact of AI-suspecting comments on posts.
@@
Overall, we find that the prevalence of posts made to these art-based communities where the authors transparently self-report AI-generated use, or where commenters raise suspicion about the use of AI, is surprisingly low, affecting less than 0.5% of image-based posts in our dataset. We find that the production of transparent AI-based image posts started increasing in early 2022, peaked in August 2022 and has decreased since, while comments suspecting the use of AI started rising only months later, and may not have declined as much. These trends were similar between communities that disallow AI and those that do not explicitly do so. Further, we identify that transparent AI posts are more likely to be made by newcomers to a community, and that newcomers who initially transparently share AI-generated images often continue to share AI-generated images, indicating that their interest in these communities has been enabled by the availability of these technologies. Finally, we find that image-based posts that are suspected of using AI have been receiving exceedingly negative comments over the period of the data collection.
@@
More broadly, our findings set the stage for understanding the prevalence, impact, and reception of AIGC in online visual art communities.
2. Background
@@ 2.1. Online Communities
Online communities have been an important object of study for internet researchers. One theory that applies to online communities, and is especially relevant to the work here, is the concept of communities of practice (CoP) (Wenger, 1999, 1996; Lave, 1991; McDonald and Cater-Steel, 2016; Wenger et al., 2002), where groups of people who share a common interest come together to explore these ideas, share, and improve their practice. CoP theory emphasizes that members engage in collective learning through shared practices, discussions, and activities, as is the case in Reddit art-based communities. Broadly speaking, online communities often wrestle with tradeoffs between different priorities, such as growth and attracting newcomers, sustainability, content quality, and fostering a sense of virtual belonging (Kairam et al., 2022; Lin et al., 2017; Hwang and Foote, 2021; Kraut et al., 2012). Recent work on Reddit mapped subreddits into five archetype, which relate to different ways they produce a sense of community (Prinster et al., 2024). In this work, we are interested in art communities, which are probably most aligned with the Content Generation archetype from that study, and aim to understand how such a community of practice might be impacted by AIGC.
@@
Indeed, the recent emergence of AIGC may present a boon for online communities, but could also put community values under threat. The Diffusion of Innovations theory, formulated by Rogers (Rogers et al., 2014; Rogers and Adhikarya, 1979), focuses on the process by which individuals adopt new innovations and ideas, suggesting that different communities may also adapt innovations in diverse ways. Weld et al. (Weld et al., 2024) identify that Reddit communities put emphasis on diverse values, and often prioritizing one comes at the cost of another – for example, the quality of content can be in direct conflict with the need to grow communities and be self-sustaining. Prior work has identified that the choice of whether to allow or disallow AI can force subreddit moderators to contend with which of these values to codify as rules (Lloyd et al., 2025b, a). These decisions can have consequences beyond individual subreddits; recent work has demonstrated that simple changes like exposing moderation decisions significantly impact the behaviors of users on the Reddit platform (Jhaver et al., 2024). As communities content with the decision of how to moderate AI-generated content, this study provides initial evidence for how these decision may shape participation in online communities.
2.2. AI-Generated Images
@@
Text-to-image models, such as DALL-E, Stable Diffusion, and Midjourney, have exhibited exceptional performance in generating both abstract and photorealistic images. Prior to the advent of these models, Generative Adversarial Networks (GAN)-based models were employed to generate images (Brock et al., 2018; Goodfellow et al., 2014; Zhang et al., 2021). With such popular image-generation tools available since 2022, the concept of AI-generated content have been increasingly impacting online communities.
@@
Since these tools are so advanced, one key challenge is the automatic detection or manual identification of AI-generated images. Efforts to determine whether an image was generated by generative models reveal the challenges humans face in accurately identifying AI-generated images (Lu et al., 2024; Nightingale and Farid, 2022), with a misclassification rate of 38.7% (Lu et al., 2024). Furthermore, even the top-performing AI-generated image detection model achieves a misclassification rate of 13% (Lu et al., 2024).
@@
A few papers have specifically examined the detection of AI-generated art (Ha et al., 2024; Wang et al., 2023a). Ha et al. (Ha et al., 2024) showed that an approach combining human experts and automatic detectors was most effective, while non-artist crowdworkers generally were unable to discern between human-created art and AI-generated images.
@@
2.3. AI-generated Art and Online Communities
@@
AI-generated art is not just as an attempt for fabrication or low-effort creation. Several papers make the case for AI-generated art creations as authentic and creative (Mazzone and Elgammal, 2019; Garcia, 2024). Garcia (Garcia, 2024) specifically investigates AI-generated art produced by advanced technologies like DALL-E. The paper provides a comprehensive analysis of the challenges associated with generative AI in art, including issues of authenticity, intellectual property, ethical concerns, the influence on traditional art practices, and market oversaturation.
@@
The discussions of challenges and opportunities in AI art extend to the context of online art communities. Communities of practice may have reasons to adopt or reject AI-generated art. For example, AI art may be seen as not contributing to improving their practice, or lacking creativity. On the other hand, AI art could be viewed as a new avenue for contribution, or a way to lower the barriers for newcomers to participate (Kraut et al., 2012).
@@
Some recent works directly explore the impact of AI-generated images on different artistic communities  (Wei and Tyson, 2024; Guo et al., 2024; Zhou and Lee, 2024). Wei and Tyson (Wei and Tyson, 2024) examined the spread of AI-generated images on Pixiv, a Japanese online community for artists, where artists are required to explicitly label their submissions as either human-generated or AI-generated. The paper reveals that the introduction of AI led to a 50% increase in new artworks, but no corresponding increase in the number of views or comments. Additionally, the introduction of AI prompted an increase in the number of new creators, but a 4.3% decrease in new creators of human-generated content. Another exploratory work (Guo et al., 2024) focused on the artist community DeviantArt which also explicitly allows AIGC. Despite DeviantArt’s requirement for users to indicate the use of AI techniques, the authors identified AI-generated artworks that did not explicitly disclose AI usage, exposing the complicated motivation of contributors. Zhou and Lee (Zhou and Lee, 2024) looked at data from an undisclosed online community focused on the creativity and productivity of artists who adopted generative AI, providing evidence for a broadening of participation and success in the community examined.
@@
The papers mentioned above examined dedicated art communities, with centralized policies and rules. In this work, we are focused on Reddit, and particularly on the decentralized, mostly-self governed (Fiesler et al., 2018) communities on this site, and especially those dedicated to art. Recent work has shown that subreddit moderators are already grappling with AIGC (Lloyd et al., 2025b, a). Subreddit moderators are creating rules that regulate AI use in their communities, clarify norms, and enforce policies. By leveraging the organizational structure of Reddit’s subreddits, we are able to analyze data based on the distinct characteristics of different communities, including their own rules — namely, if the community explicitly disallows AI, or is neutral towards it in its rules. Our analysis also covers questions such as submissions volume and number of participants, in light of the different community types.
3. Dataset
@@
To investigate online discussions within art communities regarding AI-generated images, we undertook a careful process to select data from Reddit communities (subreddits) dedicated to art. Then, to collect posts and comments related to AI-generated images, we used an initial keyword-based filtering using a list of generative AI keywords, presented in this section, and LLM-based labeling, detailed in the next section.
@@
Since there is no known existing list of art-related subreddits from prior work, we used the Reddit API combined with several external sources to construct an initial comprehensive list. These sources include art-related subreddits retrieved using the PRAW API’s search function [^1] https://praw.readthedocs.io/en/latest/index.html, a Reddit page detailing Reddit’s largest communities [^2] https://www.reddit.com/best/communities/1/, and an unofficial article describing the best art subreddits [^3] https://www.reddit.com/r/redditlists/comments/141nga/list_of_art_subreddits/. This initial step yielded a list of 276 subreddits. We further narrowed this initial list of subreddits by manually examining the description of each subreddit to ascertain its relevance to art. When descriptions were insufficient for a decision, we examined a random sample of five posts from the subreddit. A subreddit was considered to be related to art if it featured visual art, regardless of whether the art posted in the subreddit is created by the authors of the posts, and could be focused on a specific genre of art, or be more general. Subreddits dedicated to memes, videos, or nudity were excluded. This elimination step resulted in 141 relevant subreddits. Finally, since we wanted to focus on AI use and response to AI in broader art communities, we excluded the “aiArt” subreddit, resulting in 140 relevant subreddits.
@@
After identifying the relevant subreddits, we proceeded to collect corresponding posts. We utilized publicly available data hosted on Academic Torrents, which contains files for the top 40,000 subreddits, spanning from the beginning of the creation of each subreddit through the end of 2023 (Watchful1, 2024). This data is mostly building on the Pushshift API (Baumgartner et al., 2020). Each data record corresponds to the Pushshift structure [^4] https://github.com/pushshift/api, including fields such as ID, date, title, author, and attached media. Given that these torrents only cover the top 40,000 subreddits, we successfully obtained torrents for 57 subreddits from our list, resulting in a final list of 57 subreddits for our analysis (see Appendix A).
@@
With the post and comment data from the subreddit submissions, we set out to identify potential discussions about generative AI by filtering for submissions that include media items, and by using a list of AI-related keywords. First, we focused only on posts that included a media item, such as an image or video. This focus allows us to analyze discussions related to specific AI-generated art instances, rather than general discussions about art and generative AI. Second, we filtered these posts using a list of keywords likely to appear in discussions about AI-generated images. We began with an initial list of keywords and searched for them within our set of subreddits. This initial list included various keywords, starting from general phrases such as “generative art” and ending with specific text-to-image tools names. By examining the discussions in the retrieved posts, we iteratively refined, filtered, and expanded the list of keywords. Each of these keywords might be a word or a phrase, and mostly includes combinations with the phrases “AI” or “artificial intelligence”, or names of text-to-image tools for creating AI-generated images. This keyword selection process was done iteratively and purposefully to achieve a robust set, resulting in a final set of 68 keywords (see Appendix B). We then identified occurrences of these 68 keywords in the titles, bodies, and comments of the posts in our 57 subreddits. This search yielded 28,864 posts or comments containing at least one of the keywords.
@@
Our final dataset includes 28,405 text items related to AI-generated images, from 14,574 distinct posts. Of these, 6,393 (43.9%) posts feature the keywords in the title or body, while 8,181 (56.1%) posts have keywords only in the comments. The dataset includes several fields: ID, post title, post body, complete comments tree, author, timestamp, detected keywords, subreddit name, and more. In the next sections, we enhance the dataset with the classification of the texts based on their relatedness to claims about AI-generative images, as well as the categorization of the included subreddits. The dataset, with the classification labels and subreddit categories, will be made available upon publication of the paper.
4. Classification of the Dataset Using LLMs